Show that if 5/3< 2x< 11/3, then x∈{y∈R such that |y- 4/3| < 1/2}.

1
Expert's answer
2022-06-07T19:11:10-0400
"\dfrac{5}{3}<2x<\dfrac{11}{3}"

"\dfrac{5}{6}<x<\dfrac{11}{6}"

"\dfrac{5}{6}-\dfrac{4}{3}<x-\dfrac{4}{3}<\dfrac{11}{6}-\dfrac{4}{3}"

"-\dfrac{1}{2}<x-\dfrac{4}{3}<\dfrac{1}{2}"

Then


"|x-\dfrac{4}{3}|<\dfrac{1}{2}, if\ \dfrac{5}{3}<2x<\dfrac{11}{3}"